The difference of 2 numbers is 12. Twice the smaller number is equal 3 less than the larger number.
Answer by oscargut(891)

Solution:
Let x = smallest number
Let y = greatest number
2x = y-3
y-x = 12
Then:
y = x+12
2x = x+12-3
2x-x = 9
x = 9
y = 9+12 = 21
Answer: x = 9, y = 21
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com